You have been hired by Kia as manager for its Pakistan operations. Assume following is the shortrun production function at their assembly plant outside Karachi: Q = 10L2 – 0.5 L3 where L is variable input labor, Q is output of Cars assembled . Find the ranges of the three stages of production. 


1
Expert's answer
2021-01-26T08:50:10-0500
"\frac {\delta Q}{\delta L}=20L-1.5L^2"

"\frac {\delta ^2 Q}{\delta L^2}=20-3L"

"L=\frac {20}{3}"


"Q=296"